OpenStudy (anonymous):

Can someone help me work out these problems? 1. y=3x^2-7 2. y=-x^2-2 3. y=-2x^2+10x-1 4. y=x^2-3x+2

OpenStudy (anonymous):

ill help you if you help me?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

The first four terms of a sequence are shown below: f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 Which of the following functions best defines this sequence? A. f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 B. f(1) = 7, f(n + 1) = f(n) - 3; for n ≥ 1 C. f(1) = 7, f(n + 1) = f(n) - 4; for n ≥ 1 D. f(1) = 7, f(n + 1) = f(n) + 4; for n ≥ 1

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

No we haven't done that yet. Sorry!

OpenStudy (anonymous):

dang. for all 4 of your equations, just copy and paste the equations into google and a thing will pop up with what itll look like.